# A Banach *-algebra associated with the action of a group on a topological space

### v1

@@ 2.2. Representations of $\ell^{1}(\Sigma)$
We will introduce a family of representation of $\ell^{1}(\Sigma)$ that will be of help in the following section. In [Tom92, Chapter 4], given $x\in X$ and a unitary representation of the isotropy subgroup $G_{x}$ it is shown how to obtain a representation of $C^{*}(\Sigma)$. As $C^{*}(\Sigma)$ and $\ell^{1}(\Sigma)$ have the same theory of representations, we also obtain (via restriction) a representation of $\ell^{1}(\Sigma)$. We are interested in the representation, that we will denote $\pi_{x}$, obtained when the unitary representation of $G_{x}$ we start with is the trivial representation (on a Hilbert space of dimension 1, that is, the complex numbers $\mathbb{C}$). We will concretely describe the representation $\pi_{x}$ in this case.
So, let’s fix $x\in X$ and consider the isotropy subgroup $G_{x}$. We can describe the left quotient space $G/G_{x}=\{s_{\beta}G_{x}:\beta\in J_{x}\}$ for the set of representatives $\{s_{\beta}:\beta\in J_{x}\}$. We consider the Hilbert space $H_{x}$ with orthonormal basis $\{v_{\beta}:\beta\in J_{x}\}$. Then, the representation $\pi_{x}:\ell^{1}(\Sigma)\to B(H_{x})$ is given by the following:
@@
  • Given $s\in G$ and $\beta\in J_{x}$, there exists a unique $\gamma\in J_{x}$ such that $ss_{\beta}$ belongs to the coset $s_{\gamma}G_{x}$. Then, we have that $\pi_{x}(\delta^{s})v_{\beta}=v_{\gamma}$.
As $\pi_{x}$ is multiplicative (and linear), this determines the representation. Then, given an element $f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma)$, we have that
$\pi_{x}\left(\sum_{s\in G}f_{s}\delta^{s}\right)=\sum_{s\in G}\pi_{x}(f_{s})\pi_{x}(\delta^{s})\in B(H_{x}).$
@@ 3. The ideal intersection property and consequences
The objective of this section is to obtain a dictionary between dynamical properties of $\Sigma$ and analytic-algebraic properties of $\ell^{1}(\Sigma)$. The main tool in order to do that is Theorem 3.10, which will be our first objective. Before being able to prove this theorem, we will have to do some technical work. Many of the results here are heavily inspired by [dJST12], but generalized for an arbitrary countable discrete group instead of $\mathbb{Z}$.
@@
$C(X)^{\prime}=\left\{f\in\ell^{1}(\Sigma):fg=gf\text{ for every }g\in C(X)\right\}.$
It’s clear that $C(X)\subset C(X)^{\prime}$. We characterize the commutant in the following proposition. We recall that, given $g\in C(X)$, $\text{supp}\left(g\right)=\overline{\{x\in X:g(x)\neq 0\}\}}$ denotes the support of $g$.
@@
$C(X)^{\prime}=\left\{f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma):\text{supp}\left(f_{s}\right)\subset\text{Fix}(s)\text{ for every }s\in G\right\}.$
@@
Let $f=\sum_{s\in G}f_{s}\delta^{s}\in C(X)^{\prime}$. This means that, for every $g\in C(X)$, we have that $fg=gf$. We compute:
$\begin{split}fg=\left(\sum_{s\in G}f_{s}\delta^{s}\right)g=\sum_{s\in G}f_{s}\alpha_{s}(g)\delta^{s}=\sum_{s\in G}f_{s}(g\circ\sigma_{s}^{-1})\delta^{s},\end{split}$
$\begin{split}gf=g\left(\sum_{s\in G}f_{s}\delta^{s}\right)=\sum_{s\in G}gf_{s}\delta^{s}.\end{split}$
Then, we must have $f_{s}g\circ\sigma^{-1}_{s}=f_{s}g$ for every $s\in G$ and for every $g\in C(X)$. Fix $s\in G$ and let $x\in X$ such that $f_{s}(x)\neq 0.$ Then we have that $g(\sigma_{s}^{-1}(x))=g(x)$ for every $g\in C(X)$. This implies that $x=\sigma_{s}^{-1}(x)$, which is equivalent to $\sigma_{s}(x)=x$, i.e., $x$ belongs to $\text{Fix}(s)$. This means that $\text{supp}\left(f_{s}\right)\subset\text{Fix}(s)$. ∎
@@
We say that $\Sigma$ is free if $\text{Aper}(\Sigma)=X$. We say that $\Sigma$ is topologically free if $\text{Aper}(\Sigma)$ is dense in $X$,i.e., $\overline{\text{Aper}(\Sigma)}=X$.
As we noted in the introduction, the notion of topologically freeness will be key in order to work with $\ell^{1}(\Sigma)$. As the name indicates, this is a notion of the action being free in a topological sense as, if the action is topologically free, then from a topological standpoint we have that generic points have free orbit.
@@
$\text{Aper}(\Sigma)=\bigcap_{s\in G\setminus\{e\}}\text{Fix}(s)^{c}.$
$X$ is compact and Hausdorff, so it is a Baire space. As $G$ is countable, we have that $\text{Aper}(\Sigma)$ is dense if and only if $\text{Fix}(s)$ has empty interior for every $s\in G\setminus\{e\}$.
Remark 3.4.
If $X$ is finite, then every point of $X$ has an infinite stabilizer subgroup, so the action cannot be topologically free. This means that if the action is topologically free then $X$ needs to be an infinite space.
@@
$\Sigma$ is topologically free if and only if $C(X)=C(X)^{\prime}$.
@@
We just note that a nonzero continuous function has nonempty support. Then, the result follows from the above remark and Proposition 3.1. ∎
We will now prove a couple of technical lemmas that are needed to prove Theorem 3.10.
@@
Let $x\in X$ and $k\in G$ such that $x\neq\sigma_{k}(x)$. Then, there exists an open neighborhood $U$ of $x$ and a unimodular function $\theta\in C(X)$ such that $\theta$ is equal to $1$ in $U$ and
$\theta(y)\overline{\theta(\sigma_{k}^{-1}(y))}=i\in\mathbb{C}\text{ for every }y\in U.$
@@
We have that $\sigma^{{-1}}_{k}(x)\neq x$, so there exists open neighborhoods $U$ and $U^{\prime}$ disjoint and closed disjoint subsets $C$ and $C^{\prime}$ such that $x\in U$, $\sigma^{-1}_{k}(x)\in U^{\prime}$, $U\subset C$ and $U^{\prime}\subset C^{\prime}$. We can assume that $\sigma_{k}^{-1}(U)$ is included in $C^{\prime}$ by replacing $U$ with $U\cap\sigma_{k}(U^{\prime})$. Using Urysohn’s Lemma we get a function $\widetilde{\theta}\in C(X)$ such that $\widetilde{\theta}$ is equal to $1$ in $C^{\prime}$ and that it is equal to 00 in $C$. We define
$\theta=\exp\left(\frac{-i\pi\widetilde{\theta}}{2}\right)\in C(X),$
that satisfies what the lemma says. ∎
Lemma 3.7.
Let $x\in X$, and let $k_{1},\ldots,k_{N}\in G$ such that $\sigma_{k_{i}}(x)\neq x$ for every $i=1,\ldots,N$. Then, there exists $U$ an open neighborhood of $x$ and unimodular functions $\theta_{1},\ldots,\theta_{2^{N}}\in C(X)$ that verify the following property: if $f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma)$, we consider
$f^{\prime}=\frac{1}{2^{N}}\sum_{l=1}^{2^{N}}\theta_{l}f\overline{\theta_{l}};$
then:
@@
    $f_{e}^{\prime}=f_{e}$;
  2. (2)
    $f_{k_{l}}^{\prime}(y)=0$ for every $y\in U,l=1,\ldots,N;$
where the $f_{j}^{\prime}$ are from the decomposition of $f^{\prime}=\sum_{s\in G}f_{s}^{\prime}\delta^{s}$.
Proof.
Let $f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma)$ and $h\in C(X)$. We compute the following:
$\begin{split}\frac{1}{2}\left(hf\overline{h}+\overline{h}fh\right)&=\frac{1}{2}\left(h\left(\sum_{s\in G}f_{s}\delta^{s}\right)\overline{h}+\overline{h}\left(\sum_{s\in G}f_{s}\delta^{s}\right)h\right)\\ &=\frac{1}{2}\left(\sum_{s\in G}hf_{s}\alpha_{s}(\overline{h})\delta^{s}+\sum_{s\in G}\overline{h}f_{s}\alpha_{s}(h)\delta^{s}\right)\\ &=\frac{1}{2}\sum_{s\in G}f_{s}\left(h\alpha_{s}\left(\overline{h}\right)+\overline{h\alpha_{s}\left(\overline{h}\right)}\right)\delta^{s}\\ &=\frac{1}{2}\sum_{s\in G}f_{s}2\text{Re}\left(h\alpha_{s}\left(\overline{h}\right)\right)\delta^{s}=\sum_{s\in G}f_{s}\text{Re}\left(h\alpha_{s}\left(\overline{h}\right)\right)\delta^{s}.\end{split}$
So, for every $f\in\ell^{1}(\Sigma)$ and $h\in C(X)$ we have:
(1) $\frac{1}{2}\left(hf\overline{h}+\overline{h}fh\right)=\sum_{s\in G}f_{s}\text{Re}\left(h\alpha_{s}\left(\overline{h}\right)\right)\delta^{s}.$
Let us now prove the lemma. We will prove it by induction on $N$. Let $N=1$, we have $k_{1}\in G$ such that $x\neq\sigma_{k_{1}}(x)$. By the previous lemma we have an open neighborhood $U$ of $x$ and a unimodular function $\theta_{0}\in C(X)$ that is equal to $1$ in U and
$\theta_{0}(y)\overline{\theta_{0}}\left(\sigma_{k_{1}}^{{-1}}(y)\right)=i\in\mathbb{C}\text{ for every }y\in U.$
We set $\theta_{1}=\theta_{0}$ and $\theta_{2}=\overline{\theta_{0}}$. Let $f=\sum_{s\in G}f_{s}\delta^{s}\in\ell^{1}(\Sigma)$. By (1), we have that
$\begin{split}f^{\prime}=\frac{1}{2}\left(\theta_{1}f\overline{\theta_{1}}+\theta_{2}f\overline{\theta_{2}}\right)=\frac{1}{2}\left(\theta_{0}f\overline{\theta_{0}}+\overline{\theta_{0}}f\theta_{0}\right)=\sum_{s\in G}f_{s}\text{Re}\left(\theta_{0}\alpha_{s}\left(\overline{\theta_{0}}\right)\right)\delta^{s}.\end{split}$
Then, we have:
  1. (1)
    $f_{e}^{\prime}=f_{e}\text{Re}\left(\theta_{0}\alpha_{e}\left(\overline{\theta_{0}}\right)\right)=f_{e}\text{Re}\left(\theta_{0}\overline{\theta_{0}}\right)=f_{e}\text{Re}\left(|\theta_{0}|^{2}\right)=f_{e}$ because $\theta_{0}$ is unimodular;
@@
    $f_{k_{1}}^{\prime}(y)=f_{k_{1}}(y)\text{Re}\left(\theta_{0}(y)\alpha_{k_{1}}\left(\overline{\theta_{0}}(y)\right)\right)=f_{k_{1}}(y)\text{Re}\left(\theta_{0}(y)\overline{\theta_{0}}\left(\sigma_{k_{1}}^{{-1}}(y)\right)\right)=f_{k_{1}}(y)\text{Re}(i)=0.$
Now we do the inductive step. Suppose we have $\sigma_{k_{1}}(x),\ldots,\sigma_{k_{N-1}}(x)$, all distinct from $x$, and we have unimodular functions $\widetilde{\theta}_{1},\ldots,\widetilde{\theta}_{2^{N-1}}\in C(X)$ and an open set $\widetilde{U}$ of $x$ such that if $f=\sum_{s\in G}f_{s}\delta^{s}$ and we denote
$\widetilde{f}=\frac{1}{2^{N-1}}\sum_{l=1}^{2^{N-1}}\widetilde{\theta}_{l}f\overline{\widetilde{\theta}_{l}}=\sum_{s\in G}\widetilde{f}_{s}\delta^{s},$
@@
    $\widetilde{f}_{k_{l}}(y)=0$ for every $y\in\widetilde{U}$ y $l=1,\ldots,N-1.$
We have $\sigma_{k_{N}}(x)\neq x$, so by the previous lemma we have a unimodular function $\theta_{0}$ and an open neighborhood $V$ of $x$ such that $\theta_{0}$ is equal to $1$ in $V$ and
$\theta_{0}(y)\overline{\theta_{0}}\left(\sigma_{k_{N}}^{{-1}}(y)\right)=i\in\mathbb{C}\text{ for every }y\in V.$
We set $U=\widetilde{U}\cap V$. We have
$\begin{split}\frac{1}{2}\left(\theta_{0}\widetilde{f}\overline{\theta}_{0}+\overline{\theta}_{0}\widetilde{f}\theta_{0}\right)&=\frac{1}{2}\left(\theta_{0}\left[\frac{1}{2^{N-1}}\sum_{l=1}^{2^{N-1}}\widetilde{\theta}_{l}f\overline{\widetilde{\theta}_{l}}\right]\overline{\theta_{0}}+\overline{\theta_{0}}\left[\frac{1}{2^{N-1}}\sum_{l=1}^{2^{N-1}}\widetilde{\theta}_{l}f\overline{\widetilde{\theta}_{l}}\right]\theta_{0}\right)\\ &=\frac{1}{2^{N}}\left(\sum_{l=1}^{2^{N-1}}(\theta_{0}\widetilde{\theta}_{l})f\overline{(\theta_{0}\widetilde{\theta}_{l})}+\sum_{l=1}^{2^{N-1}}(\overline{\theta_{0}}\widetilde{\theta_{l}})f\overline{(\overline{\theta_{0}}\widetilde{\theta_{l}})}\right);\\ \end{split}$
setting $\theta_{l}=\theta_{0}\widetilde{\theta}_{l}$ for $l=1,\ldots,2^{N-1}$ and $\theta_{l}=\overline{\theta_{0}}\widetilde{\theta}_{l-2^{N-1}}$ for $l=2^{N-1}+1,\ldots,2^{N}$, we get from the above equation
@@
Now, using (1), we have that
$\begin{split}f^{\prime}=\frac{1}{2}\left(\theta_{0}\widetilde{f}\overline{\theta}_{0}+\overline{\theta}_{0}\widetilde{f}\theta_{0}\right)=\sum_{s\in G}\widetilde{f_{s}}\text{Re}\left(\theta_{0}\alpha_{s}\left(\overline{\theta_{0}}\right)\right)\delta^{s}.\end{split}$
Finally, writing $f^{\prime}=\sum_{s\in G}f_{s}^{\prime}\delta^{s}$, we obtain:
@@
We will use the following theorem, which can be seen in [Sak71, Theorem 1.2.4].
Theorem 3.8.

### v2

@@ 2.2. Representations of $\boldsymbol{\ell^{1}(G\curvearrowright X)}$
We introduce a family of representations of $\ell^{1}(G\curvearrowright X)$ that is going to be of help in the following section. In [Tom92, Chapter 4], given $x\in X$ and a unitary representation of the isotropy subgroup $G_{x}$, it is shown how to obtain a representation of the universal crossed product $C^{*}(G\curvearrowright X)$. As $C^{*}(G\curvearrowright X)$ and $\ell^{1}(G\curvearrowright X)$ have the same theory of representations, we also obtain (via restriction) a representation of $\ell^{1}(G\curvearrowright X)$. We are interested in the representation, that we denote by $\pi_{x}$, obtained when the unitary representation of $G_{x}$ we start with is the trivial representation (on a Hilbert space of dimension 1, that is, the complex numbers $\mathbb{C}$). We concretely describe the representation $\pi_{x}$ in this case.
So, let’s fix $x\in X$ and consider the isotropy subgroup $G_{x}$. We can describe the left quotient space $G/G_{x}=\{s_{\beta}G_{x}:\beta\in J_{x}\}$ for the set of representatives $\{s_{\beta}:\beta\in J_{x}\}$. We consider the Hilbert space $H_{x}$ with orthonormal basis $\{v_{\beta}:\beta\in J_{x}\}$. Then, the representation $\pi_{x}:\ell^{1}(G\curvearrowright X)\to B(H_{x})$ is given by the following:
@@
  • Given $s\in G$ and $\beta\in J_{x}$, there exists a unique $\gamma\in J_{x}$ such that $ss_{\beta}$ belongs to the coset $s_{\gamma}G_{x}$. Then, we have that $\pi_{x}(\delta_{s})v_{\beta}=v_{\gamma}$.
As $\pi_{x}$ is multiplicative (and linear), this determines the representation. Thus, given an element $f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X)$, we have that
$\pi_{x}\left(\sum_{s\in G}f_{s}\delta_{s}\right)=\sum_{s\in G}\pi_{x}(f_{s})\pi_{x}(\delta_{s})\in B(H_{x}).$
@@ 3. The ideal intersection property and consequences
The objective of this section is to obtain a dictionary between dynamical properties of the action $G\curvearrowright X$ and analytic-algebraic properties of $\ell^{1}(G\curvearrowright X)$. The main tool in order to do this is Theorem 3.7, which is our first objective. Before being able to prove this theorem, we have to do some technical work. Many of the results here are heavily inspired by [dJST12], but generalized for an arbitrary countable infinite discrete group instead of $\mathbb{Z}$.
@@
$C(X)^{\prime}=\left\{f\in\ell^{1}(G\curvearrowright X):fg=gf\text{ for every }g\in C(X)\right\}.$
It’s clear that $C(X)\subset C(X)^{\prime}$. We characterize the commutant in the following proposition. We recall that, given $g\in C(X)$, by $\text{supp}\left(g\right)=\overline{\{x\in X:g(x)\neq 0\}}$ we denote the support of $g$.
@@
$C(X)^{\prime}=\left\{f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X):\text{supp}\left(f_{s}\right)\subset\text{Fix}(s)\text{ for every }s\in G\right\}.$
@@
Let $f=\sum_{s\in G}f_{s}\delta_{s}\in C(X)^{\prime}$. This means that, for every $g\in C(X)$, we have that $fg=gf$. We compute directly that $fg=\sum_{s\in G}f_{s}\alpha_{s}(g)\delta_{s}$ and that $gf=\sum_{s\in G}gf_{s}\delta_{s}$. As $\alpha_{s}(g)=g\circ\sigma_{s}^{-1}$, we must have $f_{s}(g\circ\sigma^{-1}_{s})=f_{s}g$ for every $s\in G$ and for every $g\in C(X)$. Fix $s\in G$ and let $x\in X$ such that $f_{s}(x)\neq 0.$ Then we have that $g(\sigma_{s}^{-1}(x))=g(x)$ for every $g\in C(X)$. This implies that $x=\sigma_{s}^{-1}(x)$, which is equivalent to $\sigma_{s}(x)=x$, i.e., $x$ belongs to $\text{Fix}(s)$. This means that $\text{supp}\left(f_{s}\right)\subset\text{Fix}(s)$. ∎
@@
We say that the action $G\curvearrowright X$ is free if $\text{Aper}(G\curvearrowright X)=X$. We say that the action $G\curvearrowright X$ is topologically free if $\text{Aper}(G\curvearrowright X)$ is dense in $X$, i.e., if $\overline{\text{Aper}(G\curvearrowright X)}=X$.
@@
$\text{Aper}(G\curvearrowright X)=\bigcap_{s\in G\setminus\{e\}}\text{Fix}(s)^{c}.$
As $X$ is compact and Hausdorff, it is a Baire space. It follows that, as $G$ is countable, we have that $\text{Aper}(G\curvearrowright X)$ is dense if and only if $\text{Fix}(s)$ has empty interior for every $s\in G\setminus\{e\}$.
@@
The action $G\curvearrowright X$ is topologically free if and only if $C(X)$ is equal to $C(X)^{\prime}$.
@@
We just note that a nonzero continuous function has support with nonempty interior. Then, the result follows from the above remark and Proposition 3.1. ∎
We now prove a couple of technical lemmas that we need in order to prove Theorem 3.7.
@@
Let $x\in X$, and let $k_{1},\ldots,k_{N}\in G$ such that $\sigma_{k_{i}}(x)\neq x$ for every $i=1,\ldots,N$. Then, there exists $U$ an open neighborhood of $x$ and unimodular functions $\theta_{1},\ldots,\theta_{2^{N}}\in C(X)$ that satisfy the following property: if $f=\sum_{s\in G}f_{s}\delta_{s}$ is an element of $\ell^{1}(G\curvearrowright X)$, and we consider
$f^{\prime}=\frac{1}{2^{N}}\sum_{l=1}^{2^{N}}\theta_{l}f\overline{\theta_{l}},$
then, if we write $f^{\prime}=\sum_{s\in G}f_{s}^{\prime}\delta_{s}$, it verifies that:
  1. (1)
    $f_{e}^{\prime}=f_{e}$,
  2. (2)
    $f_{k_{l}}^{\prime}(y)=0$ for every $y\in U$ and every $l=1,\ldots,N.$
@@
Let $f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X)$ and let $h\in C(X)$. A direct computation shows that
(1) $\frac{1}{2}\left(hf\overline{h}+\overline{h}fh\right)=\sum_{s\in G}f_{s}\text{Re}\left(h\alpha_{s}\left(\overline{h}\right)\right)\delta_{s}.$
Let us now prove the lemma. We prove it by induction on $N$. Let $N=1$, we have $k_{1}\in G$ such that $x\neq\sigma_{k_{1}}(x)$. We can take an open neighborhood $U$ of $X$ and an unimodular function $\theta_{0}$ such that $\theta_{0}$ is equal to $1$ in $U$ and equal to i-i in $\sigma_{k_{1}}^{-1}(U)$. We set $\theta_{1}=\theta_{0}$ and $\theta_{2}=\overline{\theta_{0}}$. Take $f=\sum_{s\in G}f_{s}\delta_{s}\in\ell^{1}(G\curvearrowright X)$. By Equation (1), we have that
$\begin{split}f^{\prime}=\frac{1}{2}\left(\theta_{1}f\overline{\theta_{1}}+\theta_{2}f\overline{\theta_{2}}\right)=\frac{1}{2}\left(\theta_{0}f\overline{\theta_{0}}+\overline{\theta_{0}}f\theta_{0}\right)=\sum_{s\in G}f_{s}\text{Re}\left(\theta_{0}\alpha_{s}\left(\overline{\theta_{0}}\right)\right)\delta_{s}.\end{split}$
Then, it follows:
@@
    $f_{e}^{\prime}=f_{e}\text{Re}\left(\theta_{0}\alpha_{e}\left(\overline{\theta_{0}}\right)\right)=f_{e}\text{Re}\left(\theta_{0}\overline{\theta_{0}}\right)=f_{e}$ because $\theta_{0}$ is unimodular;
@@
    $\begin{split}f_{k_{1}}^{\prime}(y)&=f_{k_{1}}(y)\text{Re}\left(\theta_{0}(y)\alpha_{k_{1}}\left(\overline{\theta_{0}}(y)\right)\right)=f_{k_{1}}(y)\text{Re}\left(\theta_{0}(y)\overline{\theta_{0}}\left(\sigma_{k_{1}}^{{-1}}(y)\right)\right)\\ &=f_{k_{1}}(y)\text{Re}(i)=0.\end{split}$
Now we do the inductive step. Suppose we have $\sigma_{k_{1}}(x),\ldots,\sigma_{k_{N-1}}(x)$, all distinct from $x$, and we have unimodular functions $\widetilde{\theta}_{1},\ldots,\widetilde{\theta}_{2^{N-1}}\in C(X)$ and an open set $\widetilde{U}$ of $x$ such that if $f=\sum_{s\in G}f_{s}\delta_{s}$ and we denote
$\widetilde{f}=\frac{1}{2^{N-1}}\sum_{l=1}^{2^{N-1}}\widetilde{\theta}_{l}f\overline{\widetilde{\theta}_{l}}=\sum_{s\in G}\widetilde{f}_{s}\delta_{s},$
@@
    $\widetilde{f}_{k_{l}}(y)=0$ for every $y\in\widetilde{U}$ and every $l=1,\ldots,N-1.$
We have $\sigma_{k_{N}}(x)\neq x$, so we can take an open neighborhood $V$ of $x$ and a unimodular function $\theta_{0}$ such that $\theta_{0}$ is equal to $1$ in $V$ and equal to i-i in $\sigma_{k_{N}}^{-1}(U)$. We set $U=\widetilde{U}\cap V$. We have
[…]
setting $\theta_{l}=\theta_{0}\widetilde{\theta}_{l}$ for every $l=1,\ldots,2^{N-1}$ and $\theta_{l}=\overline{\theta_{0}}\widetilde{\theta}_{l-2^{N-1}}$ for every $l=2^{N-1}+1,\ldots,2^{N}$, we get from the above equation that
@@
Now, using Equation (1), we have that
$\begin{split}f^{\prime}=\frac{1}{2}\left(\theta_{0}\widetilde{f}\overline{\theta}_{0}+\overline{\theta}_{0}\widetilde{f}\theta_{0}\right)=\sum_{s\in G}\widetilde{f_{s}}\text{Re}\left(\theta_{0}\alpha_{s}\left(\overline{\theta_{0}}\right)\right)\delta_{s}.\end{split}$
Finally, writing $f^{\prime}=\sum_{s\in G}f_{s}^{\prime}\delta_{s}$, we obtain:
@@
Lemma 3.6.
Suppose that $I$ is a closed ideal of $\ell^{1}(G\curvearrowright X)$ such that $I\cap C(X)=\{0\}$. Then, for every $f=\sum_{s\in G}f_{s}\delta_{s}\in I$, we have that $f_{s}$ vanishes on $\text{Aper}(G\curvearrowright X)$ for every $s\in G$.
Proof.
Let $f=\sum_{s\in G}f_{s}\delta_{s}\in I$. We want to prove that $f_{s}(x)=0$ for every $x\in\text{Aper}(G\curvearrowright X)$ and for every $s\in G$. It is enough to prove that $f_{e}(x)=0$ for every $x\in\text{Aper}(G\curvearrowright X)$ as, given that $I$ is an ideal, we have that $f\delta_{s^{-1}}$ belongs to $I$ for every $s\in G$. We fix $x\in\text{Aper}(G\curvearrowright X)$.
We consider the quotient $\ell^{1}(G\curvearrowright X)/I$. It is a Banach algebra with the following norm:
$\lVert h+I\rVert=\inf_{j\in I}\lVert h+j\rVert\text{ for every }h\in\ell^{1}(G\curvearrowright X).$
Let $q:\ell^{1}(G\curvearrowright X)\to\ell^{1}(G\curvearrowright X)/I$ be the canonical projection, which is clearly a contraction. We have that $I\cap C(X)=\{0\}$, so we can define a norm in $C(X)$ using the norm in the quotient: we set $\lVert\>\cdot\>\rVert^{\prime}=\lVert\>\cdot\>\rVert\circ q$. This norm turns $C(X)$ into a normed algebra, so by [Sak71, Theorem 1.2.4] we have that $\lVert\>\cdot\>\rVert_{\infty}\leq\lVert\>\cdot\>\rVert$$^{\prime}$. But, as $q$ is contractive, we also have that $\lVert\>\cdot\>\rVert^{\prime}\leq\lVert\>\cdot\>\rVert$$_{\infty}$, so the two norms must coincide. It follows that the restriction of $q$ to $C(X)$ is an isometry.
$\lVert f-(f_{e}+\sum_{s\in F}f_{s}\delta_{s})\rVert<\varepsilon.$
We set $b$ as $\sum_{s\in F}f_{s}\delta_{s}$ and we set $c$ as $f-(f_{e}+b)$, so we can write $f=f_{e}+b+c$, with $\lVert c\rVert<\varepsilon$. As $F$ is finite, we have $F=\{k_{1}\ldots,k_{N}\}$, and as $x$ belongs to $\text{Aper}(G\curvearrowright X)$ we have that $\sigma_{k_{i}}(x)\neq x$ for every $i=1\ldots,N$. Using Lemma 3.5 we obtain unimodular functions $\theta_{1},\ldots,\theta_{M}\in C(X)$ and an open neighborhood $U$ of $x$ such that if we write
$f^{\prime}=\frac{1}{M}\sum_{l=1}^{M}\theta_{l}f\overline{\theta_{l}}=\sum_{s\in G}f^{\prime}_{s}\delta_{s}$
then $f_{e}^{\prime}=f_{e}$ and $f_{k_{i}}^{\prime}(y)=0$ for every $y\in U$ and every $i=1,\ldots,N$. It is clear that $f^{\prime}$ belongs to $I$. We take $\varphi\in C(X)$, with $0\leq\varphi\leq 1$, supported in $U$, and such that $\varphi(x)=1$. We set $\widetilde{f}=\varphi f^{\prime}$, so we have that $\widetilde{f}$ belongs to $I$. We can write $\widetilde{f}=\sum_{s\in G}\widetilde{f}_{s}\delta_{s}$, and we have that $\widetilde{f}_{e}(x)=\varphi(x)f_{e}^{\prime}(x)=f_{e}(x)$.
Now, we have the following, where we are using that the functions $\theta_{l}$ are unimodular in order to get the first summand in the last equality:
$\begin{split}\widetilde{f}&=\varphi f^{\prime}=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}f\overline{\theta_{l}}\\ &=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}f_{e}\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}\\ &=\varphi f_{e}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}.\end{split}$
We have that $b=\sum_{s\in F}f_{s}\delta_{s}$, so using that $\varphi$ is supported in $U$ and that the functioms $f^{\prime}_{k_{i}}$ vanish in $U$ for every $i=1\ldots,N$, we obtain
$\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}b\overline{\theta_{l}}=\sum_{s\in F}\varphi f_{s}^{\prime}\delta_{s}=\sum_{i=1}^{N}\varphi f_{k_{i}}^{\prime}\delta_{k_{i}}=0.$
We have then that $\widetilde{f}=\varphi f_{e}+\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}.$ If we set $\widetilde{c}=\frac{1}{M}\sum_{l=1}^{M}\varphi\theta_{l}c\overline{\theta_{l}}$, we compute